what is the force of attraction between you and your evil twin brother located at the diametric end of earth with respect to you?
Answer #1

My weight is around 70 Kg (I know its a lil more)

So my evil brother weight is 70 kg too

Distance between us = 2r = 12.74*10^6 m

So Force of attraction = Gmm/r^2 = 6.67*10^(-11)*70*70/(12.74*10^6)^2 = 2.01*10^(-21) N (Ans) (And as it is very less i am safe from my evil twin)
Answer #2

According to gravitatinal force

F=G\frac{M^2}{d^2}

where m=70 kg is your mass and d=12740 is the mean Earth diameter so that

F=2\cdot10^{-21}N